OpenStudy (kayla0297):

[Periodic Functions and Applications] Solve sin^2x=1/4 for 0

OpenStudy (gunsodo):

From sin^2x=1/4 So sinx = 1/2 You could continue by looking up for formulas

OpenStudy (kayla0297):

Hmmm I got that far, then first solution is 30degrees, the second is 150degrees...? There must be something else I am missing, as the textbook lists two other values :S

Parth (parthkohli):

\[\sin^2 x = 1/4 \implies \sin(x) = \pm 1/2\]

Parth (parthkohli):

Remember that the square of \(-1/2\)

Parth (parthkohli):

is also \(1/4.\)

OpenStudy (kayla0297):

Ah, so while 30 and 150 degrees are the solutions when sin is positive, having sin^2x means that I must also find solutions when sin is negative?

Parth (parthkohli):

Indeed.

OpenStudy (kayla0297):

Okay, thanks for the help!

Parth (parthkohli):

What did you get for the negative values? And no problem!

OpenStudy (kayla0297):

210 and 330 degrees

Parth (parthkohli):

That is correct.

OpenStudy (anonymous):

YES